OpenStudy (anonymous):

Integrate the function. dx/(x(sqrt 9x^2-25))

OpenStudy (jamesj):

No, that doesn't work. Try this substitution, x = 5/3 sec u

OpenStudy (anonymous):

i have a question as to why it is 5/3 sec U

OpenStudy (jamesj):

because we want to deal with all of that messiness inside the square root, and this substitution will make it easy. in particular, sec because sec^2 u - 1 = tan^2 u and 5/3 because that deals with all the constants.

OpenStudy (anonymous):

are you sure its not 3/5

OpenStudy (anonymous):

because if i do it your way the answer is...... 5/3 sec^-1 (5/3x) +c

OpenStudy (jamesj):

No 5/3. We want to be able to write that as \[ \sqrt{k^2(\sec^2 u - 1)} \] and 5/3 enables us to cancel out the 9 and generate the 25

OpenStudy (anonymous):

ok so then what?

OpenStudy (jamesj):

right ... so work through it, as it seems you already substantially have.

OpenStudy (anonymous):

so its 1/5 sec^-1 (3/5x)+c?

OpenStudy (jamesj):

looks about right. arcsec is a slightly bizarre function; if you're feeling motivated, it's better to rewrite it as arctan

OpenStudy (anonymous):

ok...not too ambitious here haha

OpenStudy (anonymous):

thank you though!

OpenStudy (jamesj):

sure

OpenStudy (anonymous):

\[\Large \frac{1}{x \sqrt{9 x^2-25}}=\frac{9 x}{x^2 \sqrt{9 x^2-25}}=\\ \Large\frac{9 x}{\left(9 x^2-25\right)^{3/2} \left(\frac{5^2}{(\sqrt{9 x^2-25})^2}+1\right)}\\ \Large u=\frac{5}{\sqrt{9 x^2-25}} \] The rest is data processing.

OpenStudy (anonymous):

\[ -\frac{1}{5} \tan ^{-1}\left(\frac{5}{\sqrt{9 x^2-25}}\right) +C \] should be your answer.